OpenStudy (anonymous):

Correct Me 4 Medal !!! ^o^

OpenStudy (mwhit2101):

?

OpenStudy (anonymous):

Consider the line that passes through the point (3, -6) and has a slope of 4. Part 1: Write the equation of this line using point-slope form. (2 points) Part 2: Using your equation from part 1, rewrite this equation in slope-intercept form. Make sure to show all of your work. (2 points) Part 3: Using your equation from part 2, rewrite this equation in standard form. Make sure to show all of your work. (2 points)

OpenStudy (anonymous):

1.) y=4x-6 2.) y-6=4(x-3) y-6=4x-12 +6 +6 ---------------- [ y = 4x - 6 ] 3.) 4x + y = -6 4x + 0 = -6

OpenStudy (anonymous):

3.) 4x + y = -6 4x + 0 = -6 4x = -6 x = -1.5 4x + y = -6 0 + y = -6 y = -6

OpenStudy (jdoe0001):

how did you get Part 1?

OpenStudy (anonymous):

y-6=4(x-3) <-- by solving that :3

OpenStudy (jdoe0001):

hmm let's solve that then... \(\bf (3,-6)\qquad m=4\qquad y-y_1=m(x-x_1)\\ \quad \\ y -(-6)= 4(x - (3))\implies y+6=4x-12\implies y = 4x-18\)

OpenStudy (anonymous):

oooooohh ! ok I see what I did wrong but the others are correct ?

OpenStudy (jdoe0001):

\( \begin{array}{llll} &point-slope form&slope-intercept form\\ y -(-6)= 4(x - (3))\implies &y+6=4x-12\implies &y = 4x-18 \end{array}\)

OpenStudy (jdoe0001):

standard form would be 4x-y-18 = 0 more or less

OpenStudy (jdoe0001):

there are several acceptable standard forms... so 4x-y = 18 may also work http://www.mathsisfun.com/algebra/standard-form.html